OpenStudy (anonymous):

Solve for x: |3x + 2| = 14

OpenStudy (anonymous):

You will have two equations 3x+2 = 14 and 3x+2 = -14 That's the definition of modulus function. ! solve for x

OpenStudy (calculusxy):

The x is 4 because of 3x4=12 and 12+2=14.

OpenStudy (anonymous):

x is -16/3 as well ! there are two solutions

OpenStudy (calculusxy):

So this is a quadratic equation because it can have many values since x is the variable.

OpenStudy (anonymous):

If x>-2/3 you solve 3x+2=14 X<-2/3 you solve 3x+2=-14